40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 28 02:43:49 2013, yuta, Update, LSC, some qualitative evidence of PRMI sideband lock PRMIsideband.png
    Reply  Thu Feb 28 03:25:27 2013, yuta, Update, LSC, PR gain ~ 25 from PRMI carrier lock PRMIcarrier.png
       Reply  Thu Feb 28 03:41:31 2013, Koji, Update, LSC, PR gain ~ 25 from PRMI carrier lock 
       Reply  Sat Mar 2 05:53:15 2013, yuta, Update, LSC, stable lock of PRMI LSCMICHOLTF_PRMI.pngLSCPRCLOLTF_PRMI.pngPRMtoAS55REFL55.png
          Reply  Sat Mar 2 14:52:02 2013, rana, Update, LSC, stable lock of PRMI 
             Reply  Sun Mar 3 22:16:46 2013, Jenne, Update, LSC, LSC whitening triggering started 
                Reply  Mon Mar 4 09:55:33 2013, rana, Update, LSC, LSC whitening triggering started 
                Reply  Tue Mar 5 18:36:27 2013, Jenne, Update, LSC, LSC whitening triggering started 
Message ID: 8213     Entry time: Sat Mar 2 14:52:02 2013     In reply to: 8212     Reply to this: 8215
Author: rana 
Type: Update 
Category: LSC 
Subject: stable lock of PRMI 

  Whereas the sensing matrix coefficients for ITMX/ITMY in REFL_I indicate an actuation imbalance, the disparity in the ITMX/ITMY to AS_Q elements does not. However, they do indicate why there is a PRM to AS_Q coupling at all.

I would recommend setting up the triggering so that the REFL & AS whitening is turned on AFTER lock acquisition and using a frequency of ~100-300 Hz for the sensing matrix measurement to fix these issues.

ELOG V3.1.3-